# Mamba State-Space Models Can Be Strong Downstream Learners

### v1

@@ 5 Regularized fine-tuning using LoRA in the MambaBlock
Thus, the overall design of the MambaBlock’s hardware optimizations may be leveraged to both efficiently learn the parameter-space for the majority of time-varying parameters (thus achieving PEFT) and regularize parameters during training (thus improving fine-tuning generalization).
@@ 6 Related Work
Several recent works [42, 19, 29, 37] have studied Mamba’s ability to perform ICL. However, none of these have extensively studied Mamba’s ICL capabilities either on standard NLP benchmarks or on pure MambaBlock foundation models. In particular, foundational Mamba models’ ICL abilities were tested in [42] to learn simple function classes (e.g., logistic regression and decision trees [17]) and in  [19] to learn non-standard NLP benchmarks (i.e., task vectors [25]). While [42, 19] report Mamba’s ICL abilities rival SOTA Transformers, their utilized benchmarks were proposed as supplemental ICL studies after Transformer LLMs’ success on standard NLP benchmarks [8]. Indeed, direct evaluation of Mamba foundation models on standard NLP benchmarks does not lead to higher gains over zero-shot performance relative to comparable Transformer LLMs (demonstrated in Table 1).
@@
As in  [22], the majority of subsequent Mamba works have focused on pretraining MambaBlocks using full precision [62, 59, 1, 37]. Notably, the official implementation of Jamba [37], the Transformer-Mamba hybrid, supports mixed- and 8-bit precision, but avoids MambaBlocks when applying such quantization [31]. Similarly, the official Mamba sources advise using full precision within the MambaBlock [28, 21], cautioning against using mixed-precision due to potential recurrent sensitivities. To the best of our knowledge, no existing works have either theoretically explored the effects small input changes (e.g., due to mixed-precision) have on Mamba’s recurrent dynamics, empirically explored such effects downstream impact on fine-tuning and inference, or explored pure Mamba networks fine-tuning abilities relative to Transformer LLMs.
@@ 7 Experiments
To demonstrate the implications of Theorem 1, we explore the performance difference between running inference with full-precision pretrained weights and using mixed-precision (FP16 and BF16) weights. Model performance is measured as percent accuracy using the MMLU [26] dataset. The difference in model performance is reported as the mean divergence (i.e., absolute difference) between the original full-precision and respective mixed-precision model, averaged over {0, 1, 3, 5}-shot percent accuracy. Thus, a divergence greater than one denotes an average difference greater than one entire percentage of accuracy.
@@
Mamba pretrained checkpoints are compared to pretrained Transformer models of similar parameter counts and no more than $\sim$300B total pretraining tokens (Pythia [5], OLMo [20] 336B-token checkpoint, and Phi 1.5 [36]). We note that Pythia and Mamba models were both pretrained using the same corpus [15], allowing the fairest comparison between SSMs and Transformers. To limit extraneous numerical effects within experiments (e.g., due to parameter aggregation across multiple GPUs), all models were run using a single GPU (Nvidia A10G, 24 GB total memory). All models were evaluated using the LM evaluation harness from Eleuther AI [16]. Further experimental details are available in Appendix C. The results are available in Table 2.
From Table 2, inferencing in Pythia using FP16 and BF16 result in an average 0.13 and 0.41 full-precision divergence, respectively. Mamba displays similar averages in comparison: inferencing in Mamba using FP16 and BF16 result in an average 0.10 and 0.48 divergence, respectively. Interestingly, both SSM and Transformer architectures exhibit large divergence spikes–i.e., mean divergence greater than a percentage point–when using BF16, which occurs once for Mamba and Phi 1.5 models and twice for Pythia models. In the following, we show that such spikes may be mitigated for Mamba SSMs by combining mixed-precision with parameter-efficient adapters during fine-tuning.
7.1 Non-divergent Mamba fine-tuning
@@
Each fine-tuning run occurred on a single A10G GPU. To further limit extraneous numerical effects, the same batch size is used for all FP32, FP16, and BF16 experiments for a given model size. While this leads to hardware underutilization (i.e., non-saturated GPU memory for mixed-precision and LoRA experiments), this is necessary to guarantee no divergence is due to differences in parameter update schedules. For comparison, Pythia 160M, 410M, and 1B models are fine-tuned using the same experimental setup. The training recipe for all models was adapted from [50], with the AdamW_torch optimizer and a cosine annealing schedule. Further experimental details are available in Appendix C.
Figure 1: Mean full-precision (FP32) divergence in MMLU performance for Mamba and Pythia models. Models are fine-tuned over the Alpaca dataset [48] using different combinations of MPFT and PEFT. Full fine-tuning (i.e., no PEFT adapters) is denoted as Full.
For each Mamba and Pythia model, Figure 1 shows the mean divergence calculated between the respective FP32 Full and mixed-precision ALL/SLL LoRA fine-tuned models, averaged over {0, 1, 3, 5}-shot MMLU accuracy. Across mixed-precisions and adapter settings, Mamba displays comparable divergences to Pythia models. E.g., for FP16, Mamba demonstrates an average divergence of 0.1, compared to 0.14 for Pythia. Similarly, for BF16, Mamba demonstrates an average divergence of 0.18, compared to 0.28 for Pythia. Importantly, Mamba models do not exhibit large deviation spikes after fine-tuning (in contrast to Pythia models).
7.2 Hardware throughput and memory-utilization improvements given PEFT and MPFT
With comparable divergences to Transformers and stable dynamics, we show that MPFT and PEFT may be used to significantly increase GPU-training throughput for Mamba SSMs. To demonstrate such improvements, we utilize the previous fine-tuning settings for the Alpaca dataset. However, we now adjust the batch size to maximize throughput per MPFT and PEFT configuration.
@@ 7.2 Hardware throughput and memory-utilization improvements given PEFT and MPFT
Figure 2: Timing and memory usage calculated Mamba model-sizes and PEFT combinations. Each model was trained using the Alpaca dataset [48] dataset for three epochs and maximum sequence length 512. For each PEFT combination, the batch size was tuned to maximize GPU occupancy.
Both throughput and memory utilization improve as the number of Mamba parameters increases in Figure 2(b). Compared to the full-precision full fine-tuning of Mamba 790M (the largest model supported by an A10G’s memory capacity), evaluated MPFT and PEFT combinations result in an average 2.15 times more training tokens-per-second while reducing per-token memory utilization by an average 62.7%. Across all model sizes, evaluated MPFT and PEFT combinations result in an average 1.74 times more training tokens-per-second while reducing per-token memory utilization by an average 47.2% compared to respective full-precision fine-tuned runs.
7.3 Fine-tuning narrows the ICL gap between Mamba and Transformers
We next explore how MPFT and PEFT affect Mamba ICL performance. All Mamba pretrained models are instruction fine-tuned using ALL LoRA and the OpenHermes dataset [49] (which consists of 242,000 supervised samples). We use the training recipe of  [50], which includes BF16 utilization. Performance is evaluated using the datasets from Table 1–HellaSwag [61], PIQA [6], Arc-E [9], Arc-C [9], and WinoGrande [46]–and report the average improvement percentage of {1, 3, 5}-shot versus 0-shot (AIPSS). For comparison, Pythia pretrained models are instruction fine-tuned using the same training recipe and ALL LoRA (i.e., all Pythia linear layers are adapted)..
Figure 3 displays AIPSS for pretrained and instruction fine-tuned Mamba and Pythia models. As previously noted, pretrained Mamba models do not display similar ICL ability as comparable Pythia models on the evaluated standard NLP benchmarks. In particular, Mamba 2.8B, the largest pretrained Mamba model, displays inconsistent zero-shot improvements as the number of shots increase. However, after fine-tuning, all Mamba models larger than Mamba 130M consistently improve in ICL performance as the number of shots increase. Compared to Mamba pretrained models, which are only capable of 38% of the AIPSS compared to similar pretrained Pythia models, fine-tuned ALL LoRA Mamba models are capable of 81.5% of the AIPSS compared to similarly fine-tuned Pythia models.
@@ 7.3 Fine-tuning narrows the ICL gap between Mamba and Transformers
(b) Pythia Pretrained
(c) Mamba ALL LoRA
(d) Pythia ALL LoRA
@@
Figure 3: Fine-tuning narrows the ICL gap between Mamba and Pythia. ALL LoRA models were instruction fine-tuned on the OpenHermes [49] dataset for one epoch. Performance is reported as the average improvement percentage of {1, 3, 5}-shot versus 0-shot over five standard benchmarks.
7.3.1 Fine-tuning robustness
We show that Mamba is robust to the choice of PEFT hyperparemters. We conduct an extensive hyperparameter search across the learning rate, LoRA dimension, and number of warmup steps. From the Cartesian-product of these three parameters, 150 hyperparameter configurations were sampled and used to fine-tune Mamba 370M over the Openhermes dataset. For comparison, Pythia 410M is similarly fine-tuned using the same set of 150 hyperparameter configurations.
(a) Mamba 370M
(b) Pythia 410M
Figure 4: Fine-tuning hyperparameter search for OpenHermes. Each point is a different hyperparameter configuration. SLL LoRA was used for both models. The $x$-axis is the learning rate, the $y$-axis is resulting MMLU 5-shot performance, bubble size is the LoRA dimension, and the color is the number of warmup steps $\in\{0,1\mbox{k},2\mbox{k}\}$.
The MMLU 5-shot performance for each of the 150 Mamba and Pythia fine-tuned models is displayed in Figure 4. Pythia 410M is capable of higher performance than Mamba 370M, where the average accuracy for the former and the latter are 26.5% and 24.8%, respectively. However, Mamba 370M is much more robust to the choice of hyperparameters, with a difference of 1.5% between the minimum (23.3%) and maximum (24.8%). In contrast, Pythia 410M fine-tuned models display a large performance difference of 4.7% between the minimum (22.9%) and maximum (27.6%).
8 Discussion
@@
We’ve extensively explored Mamba’s downstream learning capabilities. Using dynamical systems theory, we’ve shown that Mamba’s recurrent dynamics are robust to small input perturbations (contrary to the current understanding of Mamba’s recurrent sensitivities). We’ve extensively confirmed this result, showing that: a) Mamba inference is robust to changes due to mixed-precision, (b) Mamba inference differences due to mixed-precision align with Transformers, (c) Mamba fine-tuning is robust to changes due to mixed-precision and PEFT, and (d) differences in downstream performance for Mamba due to MPFT and PEFT can be more robust than Transformers. Using both MPFT and PEFT, we’ve shown that instruction fine-tuning Mamba SSMs greatly narrows the previously observed ICL gap, going from only 38% (post pretraining) up to 81.5% (post fine-tuning) of the ICL abilities of similar Transformers. Furthermore, we’ve shown that combining MPFT and PEFT can more than halve training time and nearly triple memory efficiency for Mamba models.
There are significant avenues for future work. In particular, adapting Mamba’s CUDA kernels to support more aggressive low-precision PEFT methods [12] would further decrease the hardware needed to train Mamba models, while providing additional speedups. Furthermore, while the largest pure Mamba model contains 2.8B parameters, the training speedups and improved memory utilization described herein may be applied to more efficiently pretrain larger pure Mamba SSMs (e.g., 7B parameters and greater), where Mamba models may better manifest emergent abilities previously displayed by Transformers (or even manifest previously unobserved abilities).
8.1 Broader Impact
The Mamba models considered are all LLMs, and thus have the same potential positive and negative societal impacts as other LLMs (e.g., hallucinations). Furthermore, fine-tuning is known to possibly erode existing LLM guardrails, and thus our methods may be adapted for this fine-tuning use case (as is the case for all PEFT and MPFT methods). However, our work improves the quality of Mamba models for downstream applications, which may be adapted for all positive LLM applications in society (e.g., personal assistants, task automation, code completion, etc.). Finally, our work decreases the computational constraints required to train and inference Mamba SSMs, which has implications for green ML (e.g., decreased CO2 emissions, positive climate change impact, etc.). 410 GPU days were used to produce the results for this paper.
9 Acknowledgements

### v2

@@ 3.2 Hardware-aware optimizations and PEFT
The proof of Theorem 3 is available in Appendix A. The specific affects of both targeting $\mathbf{W}$ for fine-tuning and Theorem 3’s impact on generalization are ablated in Appendix F.
@@ 4 Related Work
Recent work has sought to understand how to efficiently increase Mamba’s hidden-state dimension by restructuring SSM operations using tensor contractions [11], leading to Mamba-2 models. A separate line of work has sought to understand the in-context learning (ICL) capabilities of Mamba LLMs when trained from scratch for specific tasks [40, 30]. Another line of recent work has sought to understand how hybrid Mamba-Transformer models may be directly distilled from Transformer models [54]. However, to the best of our knowledge, no existing works have either theoretically explored the effects small input changes (e.g., due to mixed-precision) have on Mamba’s recurrent dynamics, empirically explored such effects downstream impact on fine-tuning and inference, or sought to understand the effects of LoRA adaptation on modules within the MambaBlock.
@@
As previously noted, recent works [40, 30] have studied Mamba’s ability to perform ICL by training Mamba models for specific tasks. Such tasks include logistic regression, decision trees, and learning other simple function classes, following the work of [17]. We emphasize that, in this set up, relatively small Mamba models–33 million and 90 million parameters for [30] and [40], respectively–are trained from scratch for every evaluated task. Indeed, [40] notes that subsequent work is necessary to understand Mamba’s ICL capabilities for language modeling using standard natural language benchmarks, as well as for larger model sizes. Thus, our study of both the pretrained and instruction tuned ICL capabilities of Mamba/Mamba-2 LLMs for natural language tasks are complimentary to previous works.
@@ 5 Experiments
To demonstrate the implications of Theorem 2, we explore the performance difference between running inference with full-precision pretrained weights and using mixed-precision (FP16 and BF16) weights. Model performance is measured as percent accuracy using the MMLU dataset [24]. The difference in model performance is reported as the mean divergence (i.e., absolute difference) between the original full-precision and respective mixed-precision model, averaged over {0, 1, 3, 5}-shot percent accuracy. Thus, a divergence greater than one denotes an average difference greater than one entire percentage of accuracy.
Mamba pretrained checkpoints are compared to pretrained Transformer models of similar parameter counts and no more than $\sim$300B total pretraining tokens (Pythia [4], OLMo [19] 336B-token checkpoint, and Phi 1.5 [33]). We note that Pythia and Mamba models were both pretrained using the same corpus [15], allowing the fairest comparison between SSMs and Transformers. To limit extraneous numerical effects within experiments (e.g., due to parameter aggregation across multiple GPUs), all models were run using a single GPU (Nvidia A10G, 24 GB total memory). All models were evaluated using the LM evaluation harness from Eleuther AI [16]. Further experimental details are available in Appendix D. The results are available in Table 1.
@@
From Table 1, inferencing in Pythia using FP16 and BF16 result in an average 0.13 and 0.41 full-precision divergence, respectively. Mamba displays similar averages in comparison: inferencing in Mamba using FP16 and BF16 result in an average 0.10 and 0.48 divergence, respectively. Interestingly, both SSM and Transformer architectures exhibit large divergence spikes–i.e., mean divergence greater than a percentage point–when using BF16, which occurs once for Mamba and Phi 1.5 models and twice for Pythia models. Due to space constraints, Mamba-2 results for the same experiment are included in Appendix Table 3. We note that, for comparable model sizes and mixed-precision, Mamba-2 models follow identical divergence trends as respective Mamba models.
In the following, we show that the observed large divergence spikes may be mitigated for Mamba SSMs by combining mixed-precision with parameter-efficient adapters during fine-tuning.
5.1 Mamba SSMs are stable learners
@@
Figure 1: Mean full-precision (FP32) divergence in MMLU performance for Mamba, Pythia, and OpenELM models. Models are fine-tuned over the Alpaca dataset [48] using different combinations of MPFT and PEFT. Full fine-tuning (i.e., no PEFT adapters) is denoted as Full.
@@
Each fine-tuning run occurred on a single A10G GPU. To further limit extraneous numerical effects, the same batch size is used for all FP32, FP16, and BF16 experiments for a given model size. While this leads to hardware underutilization (i.e., non-saturated GPU memory for mixed-precision and LoRA experiments), this is necessary to guarantee no divergence is due to differences in parameter update schedules. For comparison, two Transformer-based LLM families of similar parameter counts are fine-tuned using the same experimental setup: Pythia (sizes 160M, 410M, and 1B) and OpenELM [37] (sizes 270M and 450M). The training recipe for all models was adapted from [50], with the AdamW_torch optimizer and a cosine annealing schedule. Further experimental details are available in Appendix D.
For each Mamba, Pythia, and OpenELM model, Figure 1 shows the mean divergence calculated between the respective FP32 Full and mixed-precision ALL/SLL LoRA fine-tuned models, averaged over {0, 1, 3, 5}-shot MMLU accuracy.
Across mixed-precisions and adapter settings, Mamba displays smaller divergences than both Pythia and OpenELM models. E.g., for FP16, Mamba demonstrates an average divergence of 0.1, compared to 0.14 for Pythia and 0.54 for OpenELM. Similarly, for BF16, Mamba demonstrates an average divergence of 0.18, compared to 0.28 for Pythia and 0.33 for OpenELM. Importantly, Mamba models do not exhibit large deviation spikes after fine-tuning, in contrast to both Pythia and OpenELM models.
@@
Figure 2: Timing and memory usage calculated Mamba model-sizes and PEFT combinations. Each model was trained using the Alpaca dataset dataset for three epochs and maximum sequence length 512. For each PEFT combination, the batch size was tuned to maximize GPU occupancy. Full fine-tuning exceeds available GPU memory (24 GB) for models greater than 790 million parameters.
Hardware throughput and memory-utilization improvements. With stable dynamics and observed divergences smaller than comparable Transformers, we show that MPFT and PEFT may be used to significantly increase GPU-training throughput for Mamba SSMs. To demonstrate such improvements, we utilize the previous fine-tuning settings for the Alpaca dataset. However, we now adjust the batch size to maximize throughput per MPFT and PEFT configuration.
Both throughput and memory utilization improve as the number of Mamba parameters increases in Figure 2(b). Compared to the full-precision full fine-tuning of Mamba 790M (the largest model supported by an A10G’s memory capacity), evaluated MPFT and PEFT combinations result in an average 2.15 times more training tokens-per-second while reducing per-token memory utilization by an average 62.7%. Across all model sizes, evaluated MPFT and PEFT combinations result in an average 1.74 times more training tokens-per-second while reducing per-token memory utilization by an average 47.2% compared to respective full-precision fine-tuned runs. Furthermore, while full fine-tuning is no longer possible on a single A10G for Mamba models greater than 790 million parameters, MPFT and PEFT allow training Mamba models up to 2.8 billion parameters on GPUs with as little as 24 GB onboard memory.
5.2 Instruction tuning impact on Mamba ICL for natural language tasks
Using both MPFT and PEFT, we next explore how instruction tuning affects Mamba and Mamba-2 ICL performance on natural language tasks. All Mamba and Mamba-2 pretrained models are instruction fine-tuned using ALL LoRA and the OpenHermes dataset [49] (which consists of 242,000 supervised samples). We use the training recipe of  [50], which includes BF16 utilization.
Zero and few-shot performance is evaluated using five standard natural language benchmarks: HellaSwag [58], PIQA [5], Arc-E [8], Arc-C [8], and WinoGrande [46]. ICL performance is reported as the average improvement percentage of {1, 3, 5}-shot versus 0-shot (AIPSS). For comparison, Pythia pretrained models are instruction fine-tuned using the same training recipe and ALL LoRA (i.e., all Pythia linear layers are adapted).
@@ 5.2 Instruction tuning impact on Mamba ICL for natural language tasks
(b) Mamba-2 Pretrained
(c) Pythia Pretrained
(d) Mamba ALL LoRA
(e) Mamba2 ALL LoRA
(f) Pythia ALL LoRA
@@
Figure 3: Instruction tuning narrows the ICL gap between Mamba and Pythia, and creates a gap from Pythia to Mamba-2 models. ALL LoRA models were instruction tuned on the OpenHermes [49] dataset for one epoch. Performance is reported as the average improvement percentage of {1, 3, 5}-shot versus 0-shot over five standard natural language benchmarks.
Figure 3 displays AIPSS for pretrained and instruction fine-tuned Mamba and Pythia models. As previously noted, pretrained Mamba models do not display similar ICL ability as comparable Pythia models on the evaluated standard NLP benchmarks. In particular, Mamba 2.8B, the largest pretrained Mamba model, displays inconsistent zero-shot improvements as the number of shots increase. While pretrained Mamba-2 models display significantly better ICL ability than Mamba models, Mamba-2 models smaller than 780 million parameters struggle.
However, after instruction tuning, all Mamba models larger than Mamba 130M consistently improve in ICL performance as the number of shots increase. Similarly, the majority of Mamba-2 models larger than Mamba 130M greatly improve in ICL performance. Thus, while pretrained Mamba and Mamba-2 models are only capable of 38% and 82% (respectively) of the AIPSS compared to similar pretrained Pythia models, instruction tuned Mamba and Mamba-2 models are capable of 81.5% and 133% of the AIPSS relative to similarly fine-tuned Pythia models. We note that a significant difference between Mamba and Mamba-2 models is the larger (by a factor of four) latent dimension.
5.3 ICL is an emergent ability of instruction tuned Mamba SSMs
We next study the emergent behavior (as a function of model size) of Mamba/Mamba-2 SSMs’ ICL abilities on natural language tasks by comparing to a larger number of Transformer-based LLMs of varying sizes. We compare to OpenELM [37] (sizes 270M, 450M, and 1.1B), TinyLlama 1.1B [59], and OLMo 1.2B [19]. To limit the emergent effects on both parameter size and pretraining token counts, we did not evaluate models greater than 2.8 billion parameters and chose open-source checkpoints as close as possible to the 300 billion total pretraining tokens used for Mamba, Mamba-2, and Pythia models. Thus, pretraining token counts for OpenELM, TinyLlama, and OLMo models were 429 billion, 503 billion, and 336 billion, respectively. We note this potentially biases ICL performance in favor of the newly evaluated Transformer-based LLMs, and that direct comparisons between Mamba, Mamba-2, and Pythia are the most fair (as these three classes of models were all pretrained on the same dataset for the same number of total pretraining tokens).
We repeat the experiments from Figure 3, where we evaluate the pretrained and instruction tuned ICL capabilities of all models. To understand the critical role of parameter counts, we group all models into two classes: LLMs containing 450 million parameters or less, and LLMs containing greater than 450 million parameters. ICL performance measured by AIPSS is displayed in Figure 4.
[…]
Figure 4: Instruction tuning improves Mamba-2 ICL performance past Transformer LLMs. ALL LoRA models were instruction fine-tuned on the OpenHermes dataset for one epoch. Performance is reported as the average improvement percentage of {1, 3, 5}-shot versus 0-shot over five standard natural language benchmarks: HellaSwag, PIQA, Arc-E, Arc-C, and WinoGrande.
From Figure 4, it is clear that pretrained SSMs and Transformers of parameter counts 270 million and less display slight or detrimental ICL abilities (i.e., few-shot performance is worse than zero-shot). For models of greater than 450 million parameters, the majority of SSMs and Transformers display positive ICL abilities, with Mamba 1.4B being an outlier in terms of poor performance. With the exception of TinyLlama at 1-shot performance and Mamba-2 2.7B for 3- and 5-shot performance, the majority of other pretrained models cluster together.
Instruction tuning greatly smooths ICL performance across both parameter classes. While instruction tuned SSMs and Transformers of 160 million parameters or fewer continue to display slight or detrimental ICL abilities, all parameters of 270 million and greater show positive ICL abilities. The instruction tuned OpenELM 450M model displays particularly impressive ICL abilities, but it is difficult to determine whether it is strictly due to architecture and/or pretraining recipe, or partially due to $143\%$ more total pretraining tokens than Mamba/Mamba-2 and Pythia models. For instruction tuned models of greater than 450 million parameters, all SSMs and Transformers show positive ICL abilities, with Mamba-2 2.7B greatly outperforming all other models (both SSM and Transformer) in this class.
Thus, in terms of ICL as a function of SSM model size, while no clear trend presents itself for pretrained models, ICL appears to emerge for instruction tuned Mamba and Mamba-2 SSMs of size 370 million and greater.
6 Discussion and Future Directions
@@
Using dynamical systems theory, we’ve shown that the recurrent dynamics of Mamba SSMs are robust to small input perturbations. We’ve extensively confirmed this result, showing that: a) Mamba inference differences due to mixed-precision align with Transformers, (b) Mamba fine-tuning is significantly more robust to changes due to mixed-precision and PEFT than Transformers, and (c) combining MPFT and PEFT can more than halve training time and nearly triple memory efficiency for Mamba models. Using both MPFT and PEFT, we’ve shown that instruction tuning Mamba and Mamba-2 SSMs greatly narrows the pretraining ICL gap on natural language tasks relative to comparable Transformer LLMs. In particular, this allows Mamba-2 SSMs to greatly outperform the ICL abilities of a large number of instruction tuned, attention-based LLMs. Furthermore, complimentary to recent studies, we’ve shown that ICL for natural language tasks can be characterized as an emergent ability of Mamba and Mamba-2 models of 370 millions parameters or greater.
There are several avenues for future work. In particular, adapting Mamba’s CUDA kernels to support more aggressive low-precision PEFT methods [12] would further decrease the hardware needed to train Mamba models, while providing additional speedups and testing the limits of the derived stability results. Future studies could further explore what known emergent abilities Mamba SSMs are (and aren’t) capable of, as well as determining new Mamba-specific emergent abilities that are potentially not possible for Transformers to achieve. Furthermore, our theoretical contributions open the door for follow up studies, both in terms of extending our stability results to more general error (and adversarial) robustness results, as well as deriving new Mamba SSM-specific LoRA schemes which leverage the implicit weight tying achieved by targeting MambaBlock modules for either PEFT accuracy or efficiency improvements.
7 Acknowledgments